OpenStudy (anonymous):

Solve the system by elimination. -2x+2y+3z=0 -2x-y+z=-3 2x+3y+3z=5

OpenStudy (anonymous):

Can you help @Luigi0210

OpenStudy (luigi0210):

Sorry, I never learned how to solve this by elimination..

OpenStudy (anonymous):

know anyone who could help?

OpenStudy (luigi0210):

@phi maybe?

OpenStudy (phi):

if we just write down the coefficients: -2x+2y+3z=0 -2x-y+z=-3 2x+3y+3z=5 -2 2 3 0 -2 -1 1 -3 2 3 3 5 The idea is make the first column all zeros below the first -2 replace the second row with first row - second row can you do that ?

OpenStudy (phi):

I'll do the first step: zero out the -2 just below the first -2 to do that we do 2nd row - 1st row: -2 -1 1 -3 <-- 2nd row 2 -2 -3 0 <--- minus 1st row. Now add 0 -3 -2 -3 <-- new 2nd row the problem is now -2 2 3 0 0 -3 -2 -3 2 3 3 5 now we want to zero out the bottom 2. replace the last row with last row + 1st row what do you get ?